How to Calculate the perimeter of a 3quarter of a circle with radius 4cm


1
Expert's answer
2021-06-28T03:24:15-0400
"Perimeter=\dfrac{3}{4}(2\pi r)"

"Perimeter=\dfrac{3}{4}(2\pi (4))=6\pi(cm)"

The perimeter of a 3quarter of a circle with radius 4cm is "6\pi" cm.